# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 6.2   The $\Lambda(1405)$ and its debated two-pole structure
The structure of the $\Lambda(1405)$ has also been intensely studied in lattice QCD. Several investigations found a spectrum with a flavor-singlet state and two octet states with masses compatible with the conventional picture, i.e. with a single pole at about 1405 MeV. The $\Lambda(1405)$ was mostly interpreted as a molecular bound state of an antikaon and a nucleon, in some cases with overlap with local three-quark operators [1035, 1036, 1037, 1038, 809, 813, 894, 1039]. These masses only refer to the lowest finite-volume energy level, see the discussion below and in [896, 895]. The authors of [1040] devised an analysis strategy to obtain the two $\Lambda(1405)$ poles with sufficient precision to distinguish them; see also Ref. [1041]. Pavao et al. [1042] argued that most lattice QCD studies so far have relied on three-quark operators to generate the physical states. Hence, signals corresponding to the meson-baryon scattering states, appearing in the finite-volume effective theory calculations, have not yet been observed. Using a pion mass of 200 MeV and single-baryon and meson-baryon operators, lattice QCD computations of the coupled channel $\pi\Sigma-\bar{K}N$ scattering amplitudes revealed a two-pole structure [896, 895] consistent with the approach based on unitary chiral perturbation theory [605].
@@
Figure 6.2: (color online) Motion of poles by changing the $SU(3)_{f}$ breaking parameter $x$. In the $SU(3)_{f}$ limit ($x=0$), the two poles belong to the $SU(3)_{f}$ singlet 1 and octet 8. The blue, red, green lines denote the $\Lambda(1380)$, $\Lambda(1405)$ and $\Lambda(1680)$. The various meson-baryon thresholds are shown as black solid lines. Reproduced from Ref. [605]. With kind permission of The European Physical Journal (EPJ).
@@
$\Lambda(1405)\frac{1}{2}^{-}$ 19% c 29.7%c
$\Lambda(1520)\frac{3}{2}^{-}$ 17% 21.2%
@@
The two-pole structure of the $\Lambda(1405)$ is, however, not undisputed. Recent conventional multichannel analyses treat the $\Lambda(1405)$ as single pole [47, 477, 478, 479, 480, 481, 482, 563, 1043]. In [1043], the $\Lambda(1405)$ was observed as a flavor-singlet state. Introducing a second pole in the analysis, the fit converged and the low-mass pole became the flavor-singlet and the high-mass pole the flavor-octet state. From now on, we speak of $\Lambda(1405)$ to denote the observed structure or the single pole, while the two poles are described as $\Lambda(1380)$ and $\Lambda(1420)$.
Table 6.2 shows the octet fraction of low-mass negative-parity $\Lambda^{*}$ resonances from two determinations within the chiral unitary approach [1013, 1017] (both results are given in Ref. [1017]) and from two quark-model calculations [535, 677]. In the chiral unitary approach, the complex of $\Lambda(1380)$ and $\Lambda(1420)$ contains one full octet state. In quark models, the octet fraction in $\Lambda(1405)$ is small and comparable to $\Lambda(1520)$, so they are both predominantly singlets. The additional octet state arising from the two-pole structure creates tension with the Gell-Mann-Okubo (GMO) formula, which connects the octet $\Lambda^{*}$ and $\Sigma^{*}$ sector to the $N^{*}$ and $\Xi^{*}$ octet family via $SU(3)_{f}$. Expressed by a mismatch parameter $\delta$, it reads
@@
Table 6.3 lists the pole masses of baryon resonances and the mismatch parameter $\delta$ assuming the custom quark model assignments. The GMO formula is well satisfied for the ground states and the $\nicefrac{{3}}{{2}}^{-}$ states. It suggests that, for instance, the $\Lambda(1690)$ mass would gradually decrease to 1510 MeV when the $s$-quark mass is tuned to the $d$-quark mass. Only a few $\Xi$ resonances with known spin and parity are identified; in the $\nicefrac{{5}}{{2}}^{+}$ sector, a $\Xi(2020)$ resonance is listed in the RPP with spin $J=\nicefrac{{5}}{{2}}$ but unknown parity. Assuming $P=+1$, the GMO formula is well satisfied. No $\Xi$ state with $J^{P}=\nicefrac{{1}}{{2}}^{-}$ is known. Using the GMO formula and assuming that the lowest-mass octet state is $\Lambda(1690)$, the mass of the $\Xi$ state with $J^{P}=\nicefrac{{1}}{{2}}^{-}$ can be predicted and is close to that of its $\nicefrac{{3}}{{2}}^{-}$ spin partner. With $M_{\Lambda}=1418$ MeV as determined in Ref. [605] and $M_{\Sigma}=1674$ MeV from the RPP, the corresponding $\Xi$ mass would be 1454 MeV, which is only 140 MeV above the ground state. This is in conflict with the observation that the lowest-mass negative-parity state is typically found more than 500 MeV above the ground state. Thus, based on the GMO formula, the scenario suggested in [1015] cannot be straightforwardly integrated into the baryon spectrum. It would require a low-mass $\Sigma^{*}$ at about 1400 MeV (see Ref. [1044, 1045, 1046]) and a $N^{*}$ below 1300 MeV, both with $J^{P}=\nicefrac{{1}}{{2}}^{-}$.
Another argument comes from the following observation. The BESIII collaboration studied the reaction $\psi(3686)\to\Lambda\bar{\Sigma}^{0}\pi^{0}+c.c.$ [1047] and found a significant $\Lambda(1405)$ contribution. It was described applying a chiral dynamics model and a Flatté-like formula for the parametrization, which are both able to describe the data. The former assumes two poles whose relative strength is fixed in the model, and only one yield is given in [1047]. Now, because the $\psi(3686)$ is an $SU(3)$ singlet state and the $\overline{\Lambda}$ an $SU(3)$ octet, $\Lambda^{*}$ resonances are produced in this process because of their octet component. Table 6.4 presents a selection of branching fractions for $\psi(3686)\to\Lambda\bar{\Sigma}^{0}\pi^{0}+c.c.$ decays. The yields of $J^{P}=\nicefrac{{1}}{{2}}^{-}$ and $\nicefrac{{3}}{{2}}^{-}$ states allow one to predict their octet components. For example, the $\Lambda(1690)\nicefrac{{3}}{{2}}^{-}$ has a slightly higher production rate compared to the $\Lambda(1670)\nicefrac{{1}}{{2}}^{-}$ and thus possibly a larger octet component, although both effects are statistically insignificant. We assume that states with $J^{P}=\nicefrac{{1}}{{2}}^{-}$ are neither dynamically favored nor suppressed compared to states with $J^{P}=\nicefrac{{3}}{{2}}^{-}$. Concerning the $\Lambda(1405)$, the quark model predicts a production strength that is larger compared to the $\Lambda(1520)$ by a factor 1.1 [535] to 1.4 [677], while in the chiral unitary approach one should expect a higher yield of the two poles $\Lambda(1380)$ and $\Lambda(1420)$ compared to the flavor-singlet $\Lambda(1520)$ by a factor 4 [1013] to 6 [1017]. By contrast, in Table 6.4 the combined yield of $\Lambda(1380)$ plus $\Lambda(1420)$ is larger compared to the $\Lambda(1520)$ only by a factor 1.9$\pm$0.9. This is compatible with the one-pole hypothesis but only incompatible with the two-pole scenario by $2.3\sigma$. A higher precision in the branching ratio may help to distinguish between the two views.
@@
Decay ×${\cal B}\times 10^{-7}$
$\psi(3686)\to\bar{\Lambda}\Lambda(1690)\frac{3}{2}^{-}$+c.c. 47$\pm$9$\pm$8
$\psi(3686)\to\bar{\Lambda}\Lambda(1670)\frac{1}{2}^{-}$+c.c. 33$\pm$7$\pm$9
$\psi(3686)\to\bar{\Lambda}\Lambda(1520)\frac{3}{2}^{-}$+c.c. 20$\pm$5$\pm$5
    $\psi(3686)\to\bar{\Lambda}\Lambda(1420)\frac{1}{2}^{-}$+c.c. }}\}} 38$\pm$4$\pm$10
$\psi(3686)\to\bar{\Lambda}\Lambda(1380)\frac{1}{2}^{-}$+c.c.
@@
Can further experiments help us understand the structure of $\Lambda(1405)$? First, precise line shapes may reveal the two-pole structure of the $\Lambda(1405)$ and, indeed, there are indications for a two-pole structure in the $\Lambda(1405)$ region [239, 1003, 1006]. Second, electroproduction may help to elucidate the structure of the $\Lambda(1405)$ region. Are both states, $\Lambda(1350)$ and $\Lambda(1430)$, seen at large momentum transfer? Do they have a hard $qqq$ core? Third, the one-pole or two-pole structure of $\Lambda(1405)$ leads to different predictions for the mass spectra of $\chi_{c0,c2}$ decays to $(\Sigma\pi)(\overline{\Sigma}\pi)$. The branching ratio for $\chi_{c0}$ / $\chi_{c2}$ decays to $\Lambda(1520)\overline{\Lambda}(1520)$ is (3.1$\pm$1.2)$\cdot 10^{-4}$ / (4.7$\pm$1.5) [1048], and we can expect the branching for decays to $\Lambda(1405)\overline{\Lambda}(1405)$ to be of similar magnitude. Experimentally, at least one neutral decay mode of $\Lambda(1405)$ or $\overline{\Lambda}(1405)$ should be chosen to avoid the contamination from $\Sigma(1385)$ or $\overline{\Sigma}(1385)$. In the chiral unitary approach, the high- and low-mass parts of the $(\overline{\Sigma}\pi)$ invariant mass spectrum contain different fractions of $SU(3)$ singlet and octet, so the recoiling $(\Sigma\pi)$ mass distribution should be different for the high- and low-mass parts. The low-mass (high-mass) region of the invariant mass spectrum for $(\Sigma\pi)$ should recoil against the low-mass (high-mass) region in the invariant mass spectrum for $(\overline{\Sigma}\pi)$. If there is only one pole, the shape of the $(\overline{\Sigma}\pi)$ invariant mass spectrum should not depend on this cut.
There is overwhelming evidence from calculations based on the chiral unitary approach and from recent lattice QCD calculations for the existence of two poles in the 1350 – 1430 MeV region. Nevertheless, it seems difficult to accommodate these two states in the general baryon spectrum. A singlet state is expected but a low-mass octet state is not. There is one perhaps remote possibility: In Ref. [1031], the two poles are found at $[1.355(16)-i0.038(14)]$ GeV and $[1.430(6)-i0.023(4)]$ GeV, and the widths are compatible at $1\sigma$. In Ref. [605], the two poles are at $[1.4154-i0.1657]$ GeV and $[1.4179-i0.0157]$ GeV, and the masses are separated by just 2.5 MeV (uncertainties are not provided). The masses and widths of the two states vary significantly when different approximations are used. Should masses and widths both be compatible, the single pole expected in quark models would be represented by two compatible poles in the molecular picture. The single pole would be responsible for the attractive interaction between $\bar{K}$ and $N$ and between $\Sigma$ and $\pi$. In the chiral unitary approach, the attractive interactions lead to two poles, but they would represent only one particle.
@@ 6.3   Hybrid resonances and cascade decays
The question of the existence of hybrid baryons, referred to at the time as “hermaphrodite baryons", dates back to the early 1980s [1049, 1050, 1051]. It was argued that, analogous to hybrid mesons, hybrid baryons should also exist. In quark models, the light $u$ and $d$ quarks, which have current masses of only a few MeV, acquire constituent masses of approximately 350 MeV. Within this constituent quark framework, gluons can be treated as effective constituents, with constituent gluon masses typically ranging between 600 MeV and 1 GeV. This leads to a novel spectroscopy, characterized by the emergence of additional $SU(6)$ multiplets, and hence a greater number of baryon states than predicted by quark excitations alone.
@@
The presence of additional multiplets arising from gluonic degrees of freedom offers a potential route to identifying hybrid baryons. One strategy involves searching for an overpopulation of states with specific quantum numbers in the nucleon sector, such as $N\nicefrac{{1}}{{2}}^{+}$ and $N\nicefrac{{3}}{{2}}^{+}$. The Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ was long considered a strong candidate for the lowest-lying hybrid nucleon [1050], particularly due to difficulties in explaining its anomalously low mass as a simple radial excitation of the nucleon ground state. Additionally, its electromagnetic transition amplitude $A_{1/2}(0)$ is large and negative, in stark contrast to the large positive value predicted by non-relativistic constituent quark models at the photon point.
However, it was shown in Ref. [958] that if the Roper were a pure hybrid baryon, the helicity amplitude $A_{1/2}(Q^{2})$ would decrease rapidly in magnitude with increasing $Q^{2}$. In contrast, quark models that interpret the Roper as the first radial excitation of the nucleon [1052, 1053] predict that its helicity amplitudes become dominant over those of the $N(1535)\nicefrac{{1}}{{2}}^{-}$, $N(1520)\nicefrac{{3}}{{2}}^{-}$, and even $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ as $Q^{2}$ increases, whereas a hybrid interpretation would imply further suppression. The most recent data, shown in Fig. 5.8, rule out the hybrid interpretation of the Roper resonance. Instead, they confirm that its core is best described as a radial excitation of the nucleon’s three-quark system, with significant contributions from the surrounding meson-baryon cloud, which becomes apparent at small virtuality $Q^{2}$ of the exchanged photon. Furthermore, the scalar helicity amplitude $S_{1/2}(Q^{2})$ is predicted to be very small for a hybrid state, which contradicts what is shown in Fig. 6.6, where the ratio $S_{1/2}/A_{1/2}$ is sizeable throughout the $Q^{2}$ range covered by the data.
Another search for a hybrid Roper-like state has been studied through the decays of $J/\psi$ [1054], produced in $e^{+}e^{-}$ collisions. The hybrid baryon may be produced in the decay $J/\psi\to\bar{p}+|Gqqq\rangle$, as depicted in Fig. 6.6. However, due to the $J/\psi$ mass of 3.1 GeV, only hybrid baryons with masses below 2.1 GeV can be generated. Decays of $\psi(2S)$ would increase the accessible mass range for hybrid searches.
@@
More recent lattice QCD predictions place the masses of the lowest hybrid baryons with $J^{P}=\nicefrac{{1}}{{2}}^{+}$ above 2 GeV. The answer to the question above becomes more complex compared to the case of a possible low-mass hybrid Roper resonance near 1.5 GeV. At such high masses, individual resonances can no longer be clearly observed as enhancements in the mass spectrum, as they overlap with other states. Furthermore, hybrid states have the same quantum numbers as ordinary $qqq$ states in the quark model and may mix with them. One may hope, however, that the heavy constituent gluon suppresses strong mixing with ordinary states, as suggested by lattice QCD projections. In that case, hybrid baryons may exhibit $Q^{2}$-dependent features similar to those predicted for the low-mass hybrid Roper. If so, a possible strategy may proceed as follows: (i) Search for all $J^{P}=\nicefrac{{1}}{{2}}^{+}$ states in the relevant mass range and compare with predictions from $qqq$ systems. (ii) Measure transition amplitudes as functions of $Q^{2}$ to identify deviations from expectations based on $qqq$ and $Gqqq$ models – particularly a rapid drop in the transverse amplitude $A_{1/2}$ and near-zero scalar amplitude $S_{1/2}$. (iii) Compare the decay pattern of the hybrid candidate with that of other baryons.
@@
Figure 6.5: Diagram for hybrid baryon production through $J/\psi$ decay. Adopted from [1054].
Table 6.5: Branching ratios (in %) for decays of the $N(2100)\nicefrac{{1}}{{2}}^{+}$ [1055] (preliminary). Their sum is 107$\pm$12.
@@
17$\pm$5 9$\pm$3 4$\pm$2 $<1$
$\Sigma K$ $N\omega$ [425] $\Delta\pi$ $N\rho$
8$\pm$3 10$\pm$4 6$\pm$2 15$\pm$4
@@
1$\pm$1 5$\pm$3 9$\pm$4 23$\pm$5
zz zzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzz Figure 6.6: (color online) Mass spectra of $N^{*}$ and $\Delta^{*}$ baryons. The mass scale is made to reproduce the masses of the nucleon and of $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$. Baryons with a large hybrid content are highlighted in blue. Adapted from [810].
@@
Fig. 6.6 shows the lattice spectrum of “ordinary” $N^{*}$ baryons with spin-parity $\nicefrac{{1}}{{2}}^{+}$ and $\nicefrac{{3}}{{2}}^{+}$, $\Delta^{*}$ baryons with $\nicefrac{{5}}{{2}}^{+}$ and $\nicefrac{{7}}{{2}}^{+}$ [800], and hybrid baryons [810]. Since lattice calculations use large quark masses, resulting in baryon masses that are systematically too high, we normalized the results to the nucleon mass. Although the finite-volume spectrum does not directly translate to physical states, it suggests the existence of baryons with large hybrid components in the higher-lying spectrum. Experimentally, the known states in the $I(J^{P})=\nicefrac{{1}}{{2}}(\nicefrac{{1}}{{2}}^{+})$ channel include the $N(1440)\nicefrac{{1}}{{2}}^{+}$, $N(1710)\nicefrac{{1}}{{2}}^{+}$, $N(1880)\nicefrac{{1}}{{2}}^{+}$, and $N(2100)\nicefrac{{1}}{{2}}^{+}$. Quark models [535, 677] predict four $N^{*}$ states with these quantum numbers, with the highest-mass state belonging to the 20-plet. Therefore, the $N(2100)\nicefrac{{1}}{{2}}^{+}$ could be interpreted as a member of the 20-plet, or a member of the fourth excitation shell, or also as a hybrid baryon.

### v4

@@ 6.2  The $\Lambda(1405)$ and its debated two-pole structure
The structure of the $\Lambda(1405)$ has also been intensely studied in lattice QCD. Several investigations found a spectrum with a flavor-singlet state and two octet states with masses compatible with the conventional picture, i.e. with a single pole at about 1405 MeV. The $\Lambda(1405)$ was mostly interpreted as a molecular bound state of an antikaon and a nucleon, in some cases with overlap with local three-quark operators Melnitchouk:2002eg , Nemoto:2003ft , Burch:2006cc , Takahashi:2009ik , Edwards:2012fx , Engel:2013ig , Hall:2014uca , Hall:2016kou . These masses only refer to the lowest finite-volume energy level, see the discussion below and in BaryonScatteringBaSc:2023ori , BaryonScatteringBaSc:2023zvt . The authors of MartinezTorres:2012yi devised an analysis strategy to obtain the two $\Lambda(1405)$ poles with sufficient precision to distinguish them; see also Ref. Molina:2015uqp . Pavao et al. Pavao:2020zle argued that most lattice QCD studies so far have relied on three-quark operators to generate the physical states. Hence, signals corresponding to the meson-baryon scattering states, appearing in the finite-volume effective theory calculations, have not yet been observed. Using a pion mass of 200 MeV and single-baryon and meson-baryon operators, lattice QCD computations of the coupled channel $\pi\Sigma-\bar{K}N$ scattering amplitudes revealed a two-pole structure BaryonScatteringBaSc:2023ori , BaryonScatteringBaSc:2023zvt consistent with the approach based on unitary chiral perturbation theory Guo:2023wes .
@@
Figure 6.2: (color online) Motion of poles by changing the $SU(3)_{f}$ breaking parameter $x$. In the $SU(3)_{f}$ limit ($x=0$), the two poles belong to the $SU(3)_{f}$ singlet 1 and octet 8. The blue, red, green lines denote the $\Lambda(1380)$, $\Lambda(1405)$ and $\Lambda(1680)$. The various meson-baryon thresholds are shown as black solid lines. Reproduced from Ref. Guo:2023wes . With kind permission of The European Physical Journal (EPJ).
@@
$\Lambda(1405)\frac{1}{2}^{-}$ - - 19% c 29.7%c
$\Lambda(1520)\frac{3}{2}^{-}$ - - 17% 21.2%
@@
The two-pole structure of the $\Lambda(1405)$ is, however, not undisputed. Recent conventional multichannel analyses treat the $\Lambda(1405)$ as single pole Matveev:2019igl , Zhang:2013cua , Zhang:2013sva , Fernandez-Ramirez:2015tfa , Kamano:2014zba , Kamano:2015hxa , Sarantsev:2019xxm , Fernandez-Ramirez:2015fbq , Anisovich:2020lec . In Anisovich:2020lec , the $\Lambda(1405)$ was observed as a flavor-singlet state. Introducing a second pole in the analysis, the fit converged and the low-mass pole became the flavor-singlet and the high-mass pole the flavor-octet state. From now on, we speak of $\Lambda(1405)$ to denote the observed structure or the single pole, while the two poles are described as $\Lambda(1380)$ and $\Lambda(1420)$.
Table 6.2 shows the octet fraction of low-mass negative-parity $\Lambda^{*}$ resonances from two determinations within the chiral unitary approach Oller:2000fj , Jido:2003cb (both results are given in Ref. Jido:2003cb ) and from two quark-model calculations Isgur:1978xj , Loring:2001ky . In the chiral unitary approach, the complex of $\Lambda(1380)$ and $\Lambda(1420)$ contains one full octet state. In quark models, the octet fraction in $\Lambda(1405)$ is small and comparable to $\Lambda(1520)$, so they are both predominantly singlets. The additional octet state arising from the two-pole structure creates tension with the Gell-Mann-Okubo (GMO) formula, which connects the octet $\Lambda^{*}$ and $\Sigma^{*}$ sector to the $N^{*}$ and $\Xi^{*}$ octet family via $SU(3)_{f}$. Expressed by a mismatch parameter $\delta$, it reads
@@
Table 6.2 lists the pole masses of baryon resonances and the mismatch parameter $\delta$ assuming the custom quark model assignments. The GMO formula is well satisfied for the ground states and the $\nicefrac{{3}}{{2}}^{-}$ states. It suggests that, for instance, the $\Lambda(1690)$ mass would gradually decrease to 1510 MeV when the $s$-quark mass is tuned to the $d$-quark mass. Only a few $\Xi$ resonances with known spin and parity are identified; in the $\nicefrac{{5}}{{2}}^{+}$ sector, a $\Xi(2020)$ resonance is listed in the RPP with spin $J=\nicefrac{{5}}{{2}}$ but unknown parity. Assuming $P=+1$, the GMO formula is well satisfied. No $\Xi$ state with $J^{P}=\nicefrac{{1}}{{2}}^{-}$ is known. Using the GMO formula and assuming that the lowest-mass octet state is $\Lambda(1690)$, the mass of the $\Xi$ state with $J^{P}=\nicefrac{{1}}{{2}}^{-}$ can be predicted and is close to that of its $\nicefrac{{3}}{{2}}^{-}$ spin partner. With $M_{\Lambda}=1418$ MeV as determined in Ref. Guo:2023wes and $M_{\Sigma}=1674$ MeV from the RPP, the corresponding $\Xi$ mass would be 1454 MeV, which is only 140 MeV above the ground state. This is in conflict with the observation that the lowest-mass negative-parity state is typically found more than 500 MeV above the ground state. Thus, based on the GMO formula, the scenario suggested in Jido:2002zk cannot be straightforwardly integrated into the baryon spectrum. It would require a low-mass $\Sigma^{*}$ at about 1400 MeV (see Ref. Belle:2022ywa , Wang:2024jyk , Lin:2025pyk ) and a $N^{*}$ below 1300 MeV, both with $J^{P}=\nicefrac{{1}}{{2}}^{-}$.
[…]
@@
Decay ${\cal B}\times 10^{-7}$
[…]
$\psi(3686)\to\bar{\Lambda}\Lambda(1380)\frac{1}{2}^{-}$+$c.$
@@
[…]
There is overwhelming evidence from calculations based on the chiral unitary approach and from recent lattice QCD calculations for the existence of two poles in the 1350 – 1430 MeV region. Nevertheless, it seems difficult to accommodate these two states in the general baryon spectrum. A singlet state is expected but a low-mass octet state is not. There is one perhaps remote possibility: In Ref. Sadasivan:2022srs , the two poles are found at [1.355(16)i0.038(14)][1.355(16)-i0.038(14)] GeV and [1.430(6)i0.023(4)][1.430(6)-i0.023(4)] GeV, and the widths are compatible at $1\sigma$. In Ref. Guo:2023wes , the two poles are at [1.4154i0.1657][1.4154-i0.1657] GeV and [1.4179i0.0157][1.4179-i0.0157] GeV, and the masses are separated by just 2.5 MeV (uncertainties are not provided). The masses and widths of the two states vary significantly when different approximations are used. Should masses and widths both be compatible, the single pole expected in quark models would be represented by two compatible poles in the molecular picture. The single pole would be responsible for the attractive interaction between $\bar{K}$ and $N$ and between $\Sigma$ and $\pi$. In the chiral unitary approach, the attractive interactions lead to two poles, but they would represent only one particle.
@@ 6.3  Hybrid resonances and cascade decays
The question of the existence of hybrid baryons, referred to at the time as “hermaphrodite baryons", dates back to the early 1980s Barnes:1977hg , Barnes:1982fj , Golowich:1982kx . It was argued that, analogous to hybrid mesons, hybrid baryons should also exist. In quark models, the light $u$ and $d$ quarks, which have current masses of only a few MeV, acquire constituent masses of approximately 350 MeV. Within this constituent quark framework, gluons can be treated as effective constituents, with constituent gluon masses typically ranging between 600 MeV and 1 GeV. This leads to a novel spectroscopy, characterized by the emergence of additional $SU(6)$ multiplets, and hence a greater number of baryon states than predicted by quark excitations alone.
@@
The presence of additional multiplets arising from gluonic degrees of freedom offers a potential route to identifying hybrid baryons. One strategy involves searching for an overpopulation of states with specific quantum numbers in the nucleon sector, such as $N\nicefrac{{1}}{{2}}^{+}$ and $N\nicefrac{{3}}{{2}}^{+}$. The Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ was long considered a strong candidate for the lowest-lying hybrid nucleon Barnes:1982fj , particularly due to difficulties in explaining its anomalously low mass as a simple radial excitation of the nucleon ground state. Additionally, its electromagnetic transition amplitude $A_{1/2}(0)$ is large and negative, in stark contrast to the large positive value predicted by non-relativistic constituent quark models at the photon point.
However, it was shown in Ref. Li:1991yba that if the Roper were a pure hybrid baryon, the helicity amplitude $A_{1/2}(Q^{2})$ would decrease rapidly in magnitude with increasing $Q^{2}$. In contrast, quark models that interpret the Roper as the first radial excitation of the nucleon Abdullah:1971aq , LeYaouanc:1972ju predict that its helicity amplitudes become dominant over those of the $N(1535)\nicefrac{{1}}{{2}}^{-}$, $N(1520)\nicefrac{{3}}{{2}}^{-}$, and even $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ as $Q^{2}$ increases, whereas a hybrid interpretation would imply further suppression. The most recent data, shown in Fig. 5.8, rule out the hybrid interpretation of the Roper resonance. Instead, they confirm that its core is best described as a radial excitation of the nucleon’s three-quark system, with significant contributions from the surrounding meson-baryon cloud, which becomes apparent at small virtuality $Q^{2}$ of the exchanged photon. Furthermore, the scalar helicity amplitude $S_{1/2}(Q^{2})$ is predicted to be very small for a hybrid state, which contradicts what is shown in Fig. 6.6, where the ratio $S_{1/2}/A_{1/2}$ is sizeable throughout the $Q^{2}$ range covered by the data.
Another search for a hybrid Roper-like state has been studied through the decays of $J/\psi$ Ping:2004wz , produced in $e^{+}e^{-}$ collisions. The hybrid baryon may be produced in the decay $J/\psi\to\bar{p}+|Gqqq\rangle$, as depicted in Fig. 6.6. However, due to the $J/\psi$ mass of 3.1 GeV, only hybrid baryons with masses below 2.1 GeV can be generated. Decays of $\psi(2S)$ would increase the accessible mass range for hybrid searches.
@@
More recent lattice QCD predictions place the masses of the lowest hybrid baryons with $J^{P}=\nicefrac{{1}}{{2}}^{+}$ above 2 GeV. The answer to the question above becomes more complex compared to the case of a possible low-mass hybrid Roper resonance near 1.5 GeV. At such high masses, individual resonances can no longer be clearly observed as enhancements in the mass spectrum, as they overlap with other states. Furthermore, hybrid states have the same quantum numbers as ordinary qqqqqq states in the quark model and may mix with them. One may hope, however, that the heavy constituent gluon suppresses strong mixing with ordinary states, as suggested by lattice QCD projections. In that case, hybrid baryons may exhibit $Q^{2}$-dependent features similar to those predicted for the low-mass hybrid Roper. If so, a possible strategy may proceed as follows: (i) Search for all $J^{P}=\nicefrac{{1}}{{2}}^{+}$ states in the relevant mass range and compare with predictions from qqqqqq systems. (ii) Measure transition amplitudes as functions of $Q^{2}$ to identify deviations from expectations based on qqqqqq and $Gqqq$ models – particularly a rapid drop in the transverse amplitude $A_{1/2}$ and near-zero scalar amplitude $S_{1/2}$. (iii) Compare the decay pattern of the hybrid candidate with that of other baryons.
@@
Figure 6.5: Diagram for hybrid baryon production through $J/\psi$ decay. Adopted from Ping:2004wz .
[…]
@@
[…]
$\Sigma K$ $N\omega$ Denisenko:2016ugz $\Delta\pi$ $N\rho$
[…]
@@
[…]
zzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzz Figure 6.6: (color online) Mass spectra of $N^{*}$ and $\Delta^{*}$ baryons. The mass scale is made to reproduce the masses of the nucleon and of $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$. Baryons with a large hybrid content are highlighted in blue. Adapted from Dudek:2012ag .
@@
Fig. 6.6 shows the lattice spectrum of “ordinary” $N^{*}$ baryons with spin-parity $\nicefrac{{1}}{{2}}^{+}$ and $\nicefrac{{3}}{{2}}^{+}$, $\Delta^{*}$ baryons with $\nicefrac{{5}}{{2}}^{+}$ and $\nicefrac{{7}}{{2}}^{+}$ Edwards:2011jj , and hybrid baryons Dudek:2012ag . Since lattice calculations use large quark masses, resulting in baryon masses that are systematically too high, we normalized the results to the nucleon mass. Although the finite-volume spectrum does not directly translate to physical states, it suggests the existence of baryons with large hybrid components in the higher-lying spectrum. Experimentally, the known states in the $I(J^{P})=\nicefrac{{1}}{{2}}(\nicefrac{{1}}{{2}}^{+})$ channel include the $N(1440)\nicefrac{{1}}{{2}}^{+}$, $N(1710)\nicefrac{{1}}{{2}}^{+}$, $N(1880)\nicefrac{{1}}{{2}}^{+}$, and $N(2100)\nicefrac{{1}}{{2}}^{+}$. Quark models Isgur:1978xj , Loring:2001ky predict four $N^{*}$ states with these quantum numbers, with the highest-mass state belonging to the 20-plet. Therefore, the $N(2100)\nicefrac{{1}}{{2}}^{+}$ could be interpreted as a member of the 20-plet, or a member of the fourth excitation shell, or also as a hybrid baryon.